OpenStudy (anonymous):

Fiber-optic cables are used widely for internet wiring, data transmission, and surgeries. When light passes through a fiber-optic cable, its intensity decreases with the increase in the length of the cable. If 1500 lumens of light enters the cable, the intensity of light decreases by 3.4% per meter of the cable.

OpenStudy (anonymous):

Can this situation be represented by a linear function? Justify your answer.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

let's analyse the situation okay

OpenStudy (anonymous):

if we suppose that light crossed one meter then the intensity deacreses by 3.4% which means 1500-1500*3.4%=1449 lumens

OpenStudy (anonymous):

now the present intensity is 1449 will across an other meter again we have 1449-1449*3.4%=1399.734 lumens

OpenStudy (anonymous):

if the operation is linear then after acrossing 2 meters the intensity should be 1500-2*(1500*3.4%)=1398 lumens